# Université Paris Cité, F-75005 Paris, France

### v1

@@ 5.1 TBA and mirror TBA
If we instead quantised the system on $\mathcal{B}$ and treated $\mathcal{C}$ as the time direction then, in the thermodynamic limit, the partition function can be computed using the Bethe ansatz. This was derived in Zamolodchikov:1989cf and the extension to also compute the excited states was derived in Dorey:1996re . We will just state the results here.
@@
We can also extract the excited states from the TBA equations by analytic continuation. This was first discussed in Dorey:1996re and further details were given in Dorey:1997rb . In Dorey:1996re it was conjectured that the TBA equation should be modified to
@@
which lead to singularities in the integrand in (72). Note that there are also singularities in the integrand due to the poles in the $S$ matrix. These poles can also give rise to additional driving terms in the TBA equation (72). Solving the TBA equations with these terms added moves us between the different Virasoro representations in our theory as detailed in Dorey:1996re . We will only solve TBA equations of the form (74) which gives us excited states in the ground state representation. In the Lee-Yang model this is the $h=-1/5$ representation.
@@
The specific choice of branch cuts of the logarithms won’t matter in our analysis but they have been carefully studied in Dorey:1997rb . The excited state energy is then given by
@@ 5.2 TBA for the GGE
The constant $C$ can be computed using the results in Bazhanov:1996aq , in particular
@@
where $C_{3}$ is given in equation (4.35), $\kappa$ in (4.16) and the combination is given in (4.34) in Bazhanov:1996aq . (Note that our TBA equation (89) differs from (4.30) in Bazhanov:1996aq where the driving term is $\kappa e^{\theta}$ instead of $e^{\theta}$. This accounts for the factor $\kappa^{5}$ in (88).)
@@
The $\alpha^{0}$ term in the integral gives the vacuum eigenvalue of $I_{1}(R)$ in the $h=-\frac{1}{5}$ representation and the $\alpha$ term gives the vacuum eigenvalue of $I_{5}(R)$ for $h=-\frac{1}{5}$. This was derived in Bazhanov:1996aq . We have started with the TBA equations for a massless theory, however we could start with a massive theory and then take the massless limit. This was done in Mazzoni:2024adh and gives the same TBA equations we are studying here.
@@
It was again verified in Bazhanov:1996aq that solving these TBA equations gives the excited state eigenvalues for $I_{1}(R)$ and $I_{5}(R)$.
@@ 5.4 Asymptotic results from the TBA
${-}\frac{\alpha^{3}}{L^{13}}\int_{-\infty}^{\infty}e^{\theta}\left(\epsilon_{3}(\theta)L^{\prime}(\epsilon_{0}(\theta)){+}\epsilon_{2}(\theta)\epsilon_{1}(\theta)L^{\prime\prime}(\epsilon_{0}(\theta)){+}\frac{1}{6}\epsilon_{1}(\theta)^{3}L^{\prime\prime\prime}(\epsilon_{0}(\theta))\right)\frac{d\theta}{2\pi}{+}O(\alpha^{4})\;.$
@@ 5.6 Non-asymptotic solutions to the TBA
Based on the results for free fermion GGEs Downing:2021mfw ; Downing:2023lnp ; Downing:2023lop we conjecture that if we add these terms to the GGE then we will have the full modular transformation. This conjecture can also be extended to the case with a finite number of charges inserted as was done for free fermions in Downing:2023lnp .
@@ 6 Conclusions and Outlook
where the $\alpha_{2n-1}$ and $\beta_{2n-1}$ are given in (5961). Based on the results for the free fermion model Downing:2021mfw ; Downing:2023lop ; Downing:2023lnp we assume that the expressions (177) only match asymptotically and that as a GGE the right hand side is a formal expression that diverges.
@@
It is also worth noting that these TBA equations can be written as the same Y system that one has for the ordinary Lee-Yang model. The original derivation of Y systems from TBA equations was given in Zamolodchikov:1991et , and in Castro-Alvaredo:2022pgz Castro-Alvaredo showed that the same Y system also encodes the TBA equations for GGEs. For the case of our Lee-Yang TBA equations (178) and (180), we define
@@
As was noted in Zamolodchikov:1991et the functions $Y(\theta)$ are periodic $Y(\theta)=Y\left(\theta+\frac{5\pi i}{3}\right)$. Hence we can further define
@@
This is the same T system first derived in Bazhanov:1994ft . However our function $t(\lambda)$ also has a dependence on $\alpha$ and hence has different analytic properties to the one defined in Bazhanov:1994ft . In Bazhanov:1994ft the asymptotic expansion of $t(\lambda)$ as $\lambda\rightarrow\infty$ gave the eigenvalues of the KdV charges in the theory. It would be interesting to understand if the asymptotic expansion of our function $t(\lambda)$ as $\lambda\rightarrow\infty$ again contains the eigenvalues of higher spin conserved charges that are present in the theory represented by our transformed GGE.
While we have provided evidence for our conjecture that the spectrum for the transformed GGE is fully encoded in the TBA equations (178) and (180) we have not provided a rigorous proof of this statement. In Downing:2023lop it was proven that for free fermions the full spectrum of the transformed GGE is encoded in the TBA equations for that model. However the proof required having the explicit expressions for the GGEs and then using Poisson summation to perform the modular transformation. Here we do not have an explicit expression for the original GGE and hence cannot attempt to use the same methods.
@@
An interesting point to mention is that in a generic 2d CFT, there exist further infinite sets of commuting conserved charges that are independent to the KdV hierarchy. In particular there exist hierarchies that are related to the ZMS-Bullough-Dodd model, see for example Fring:1992pj , and can be constructed by considering certain integrable perturbations of CFTs Mathieu:1996un (in fact there are two sets of Bullough-Dodd charges which depend on the choice of the integrable perturbation). In the case of the Lee-Yang Model that we have analysed in this paper, the KdV hierarchy and the Bullough-Dodd hierarchies exactly coincide. It is then natural to ask about GGEs with Bullough-Dodd charges inserted in them in a more general setting.
There is also the BO2 hierarchy that exists for CFTs that have a $U(1)$ current. GGEs with these charges inserted have been studied in Novaes:2021vjh ; Melnikov:2018fhb . Studying their modular properties is an open question that would be interesting to explore.
@@
The modular properties of this GGE is still an open question. The first few terms in the asymptotic expansion and their modular transforms were calculated in Iles:2013jha ; Iles:2014gra . The additional charges and their thermal correlators have recently been calculated in Ashok:2024zmw . Putting these results together could allow us to find an asymptotic expression for the modular transform of (187) similar to our results in section 3. If TBA equations for the additional charges are known then we may hope to repeat the arguments of section 5 to find the full modular transform of (187).
@@ Acknowledgements
We would like to thank O. Castro-Alvaredo and A. Konechny for useful discussions during the completion of this work. We would especially like the thank G. M. T. Watts for many extensive and valuable discussions that aided in the progression of this work, reading early drafts of the paper and suggesting useful edits.
@@ Appendix A Modular forms
In this appendix we will list the relevant facts about modular forms that appear in this paper. Proofs of the following statements can be found in Zagier2008 and most of the notation will be the same.
@@ Appendix B Construction of Charges
In this appendix we will explain how to construct the charges used throughout this paper. These charges are the zero modes of quasi-primary fields on the cylinder so we will begin by explaining how we use the algorithm of Gaberdiel in Gaberdiel:1994fs to map fields from the cylinder to the plane. We will then apply this map to the case of the quasi-primary field at level 6 that is linearly independent from the quasi-primary field that gives the KdV charge $I_{5}$. Finally we will discuss the charges in the Lee-Yang minimal model and give explicit expressions for all the charges used in this work.
@@ B.1 Mapping between the cylinder and plane
The algorithm for computing the $R_{n}$ is given in Gaberdiel:1994fs , we have listed the relevant ones for our calculations in table 5. We note that for $n$ odd and greater than 1 $R_{n}=0$ so we have not listed them in table 5. (For a general conformal map $z=f(w)$ the $R_{n}$ will be functions of $w$ and $z^{L_{0}}$ becomes $f^{\prime}(z)$.) Although we have an infinite product and the exponentials also contain infinite products these expressions can truncate to a finite one since any operator of the form $L_{n_{1}}\dots L_{n_{i}}$ with $n_{1}+\dots+n_{i}>N$ will annihilate $|\phi\rangle$.
@@ B.2 Example of a Weight 6 Field
The brackets denote normal ordering as defined in DiFrancesco:1997nk . This then gives use the field $\phi^{\text{cyl}}(w)$ on the cylinder
@@ B.3 Charges in the Lee-Yang Model
In this section we will present the charges relevant to the Lee-Yang model. These will include the KdV charges which have been calculated previously (see for example Bazhanov:1994ft ). However we will find new simpler expressions for them by using a bases of states in the Lee-Yang theory which already has null states removed. We will also calculate the charges associated with the other quasi-primary fields in the theory. While these don’t commute with the KdV charges they still appear when we take the modular transform of a GGE as see in section 4.
There is natural basis of states in the vacuum module which avoids the null vectors in the Lee-Yang model Bajnok:2014fca , that is the vacuum module is given by
@@
Hence we have no quasi-primary states at level 4. This means we have no $I_{3}$ KdV charge as has previously been pointed out in Maloney:2018hdg .
@@
This is proportional to the state which gives the KdV charge $I_{5}$. However if we map this state to the cylinder then we get an expression for the zero mode which only contains quadratic and linear terms in the Virasoro modes rather than the usual expression for $I_{5}$ which contains terms that are cubic (see for example the expression in Bazhanov:1994ft ). Hence using the representation (212) for the vacuum module leads to simpler expressions for the charges.
@@
The $I_{2n-1}$ are the KdV charges coming from a weight $2n$ quasi-primary field. They can be uniquely fixed (up to a factor) by imposing that they all commute Bazhanov:1994ft . $J_{9}$ is the zero mode of the unique quasi-primary field at level 10, it does not commute with the KdV charges. $J_{13}$ is the zero mode of the quasi-primary field at level 14 that is linearly independent to the field that gives the KdV charge $I_{13}$, it also doesn’t commute with the KdV charges.
@@ Appendix C Eigenvalues and Thermal Correlation Functions of Charges
In this appendix we explain how to compute the thermal correlation functions of the charges using the techniques of Maloney:2018hdg . These thermal correlation functions are given by modular differential operators acting on the characters of the theory.
@@ C.2 Thermal Correlation Functions
Using the method presented in the previous section to calculate sums of eigenvalues of operators, we can find the expression of thermal correlation functions in terms of modular differential operators acting on the characters of the CFT as was done in Maloney:2018hdg .
@@
From Maloney:2018hdg we know that the thermal correlator must be a weight 6 modular differential operator acting on the characters. We only have a linear $L_{0}$ term in (211), hence we must have a first order differential operator. Using the definition of the Serre derivative $D$ and Eisenstein series $E_{2k}$ in appendix A the only first order weight 6 differential operator we can construct is
@@
If we restrict to the Lee-Yang model we can find simpler expressions for the correlation functions compared to those given in Maloney:2018hdg for generic CFTs. This is because the characters satisfy a second order differential equation which can be used to reduce the order of differential operator acting on the characters.
We will demonstrate this for the case of the correlator $\langle{I_{5}}\rangle$ in the Lee-Yang model. (For the rest of this section we will drop the $i$ subscript since the specific representation won’t be important.) From Maloney:2018hdg , $\langle{I_{5}}\rangle$ in the Lee-Yang model is
@@
However we also have the modular differential equation satisfied by the two characters of Lee-Yang Gaberdiel:2008pr
@@ Appendix D Numerical algorithm for TBA
In this appendix we will briefly outline our approach to solving the TBA equations numerically. We do this by discretising the integrals into finite sums and then setting up iteration schemes. Our iteration scheme for finding the pseudo energy $\epsilon(\theta)$ is the same as the one used in equations (2.2) of Dorey:1996re with $a=1$. We will also explicitly give the iteration scheme we used to solve for the poles $\eta$ in the excited states.

### v2

@@ 5.1 TBA and mirror TBA
If we instead quantised the system on $\mathcal{B}$ and treated $\mathcal{C}$ as the time direction then, in the thermodynamic limit, the partition function can be computed using the Bethe ansatz. This was derived in [26] and the extension to also compute the excited states was derived in [27]. We will just state the results here.
@@
We can also extract the excited states from the TBA equations by analytic continuation. This was first discussed in [27] and further details were given in [28]. In [27] it was conjectured that the TBA equation should be modified to
@@
which lead to singularities in the integrand in (72). Note that there are also singularities in the integrand due to the poles in the $S$ matrix. These poles can also give rise to additional driving terms in the TBA equation (72). Solving the TBA equations with these terms added moves us between the different Virasoro representations in our theory as detailed in [27]. We will only solve TBA equations of the form (74) which gives us excited states in the ground state representation. In the Lee-Yang model this is the $h=-1/5$ representation.
@@
The specific choice of branch cuts of the logarithms won’t matter in our analysis but they have been carefully studied in [28]. The excited state energy is then given by
@@ 5.2 TBA for the GGE
The constant $C$ can be computed using the results in [18], in particular
@@
where $C_{3}$ is given in equation (4.35), $\kappa$ in (4.16) and the combination is given in (4.34) in [18]. (Note that our TBA equation (89) differs from (4.30) in [18] where the driving term is $\kappa e^{\theta}$ instead of $e^{\theta}$. This accounts for the factor $\kappa^{5}$ in (88).)
@@
The $\alpha^{0}$ term in the integral gives the vacuum eigenvalue of $I_{1}(R)$ in the $h=-\frac{1}{5}$ representation and the $\alpha$ term gives the vacuum eigenvalue of $I_{5}(R)$ for $h=-\frac{1}{5}$. This was derived in [18]. We have started with the TBA equations for a massless theory, however we could start with a massive theory and then take the massless limit. This was done in [29] and gives the same TBA equations we are studying here.
@@
It was again verified in [18] that solving these TBA equations gives the excited state eigenvalues for $I_{1}(R)$ and $I_{5}(R)$.
@@ 5.4 Asymptotic results from the TBA
${-}\frac{\alpha^{3}}{L^{13}}\int_{-\infty}^{\infty}e^{\theta}\left(\epsilon_{3}(\theta)L^{\prime}(\epsilon_{0}(\theta)){+}\epsilon_{2}(\theta)\epsilon_{1}(\theta)L^{\prime\prime}(\epsilon_{0}(\theta)){+}\frac{1}{6}\epsilon_{1}(\theta)^{3}L^{\prime\prime\prime}(\epsilon_{0}(\theta))\right)\frac{d\theta}{2\pi}$
${+}O(\alpha^{4})\;.$
@@ 5.6 Non-asymptotic solutions to the TBA
Based on the results for free fermion GGEs [15, 16, 17] we conjecture that if we add these terms to the GGE then we will have the full modular transformation. This conjecture can also be extended to the case with a finite number of charges inserted as was done for free fermions in [16].
@@ 6 Conclusions and Outlook
where the $\alpha_{2n-1}$ and $\beta_{2n-1}$ are given in (5961). Based on the results for the free fermion model [15, 17, 16] we assume that the expressions (177) only match asymptotically and that as a GGE the right hand side is a formal expression that diverges.
@@
It is also worth noting that these TBA equations can be written as the same Y system that one has for the ordinary Lee-Yang model. The original derivation of Y systems from TBA equations was given in [30], and in [31] Castro-Alvaredo showed that the same Y system also encodes the TBA equations for GGEs. For the case of our Lee-Yang TBA equations (178) and (180), we define
@@
As was noted in [30] the functions $Y(\theta)$ are periodic $Y(\theta)=Y\left(\theta+\frac{5\pi i}{3}\right)$. Hence we can further define
@@
This is the same T system first derived in [3]. However our function $t(\lambda)$ also has a dependence on $\alpha$ and hence has different analytic properties to the one defined in [3]. In [3] the asymptotic expansion of $t(\lambda)$ as $\lambda\rightarrow\infty$ gave the eigenvalues of the KdV charges in the theory. It would be interesting to understand if the asymptotic expansion of our function $t(\lambda)$ as $\lambda\rightarrow\infty$ again contains the eigenvalues of higher spin conserved charges that are present in the theory represented by our transformed GGE.
While we have provided evidence for our conjecture that the spectrum for the transformed GGE is fully encoded in the TBA equations (178) and (180) we have not provided a rigorous proof of this statement. In [17] it was proven that for free fermions the full spectrum of the transformed GGE is encoded in the TBA equations for that model. However the proof required having the explicit expressions for the GGEs and then using Poisson summation to perform the modular transformation. Here we do not have an explicit expression for the original GGE and hence cannot attempt to use the same methods.
@@
An interesting point to mention is that in a generic 2d CFT, there exist further infinite sets of commuting conserved charges that are independent to the KdV hierarchy. In particular there exist hierarchies that are related to the ZMS-Bullough-Dodd model, see for example [32], and can be constructed by considering certain integrable perturbations of CFTs [33] (in fact there are two sets of Bullough-Dodd charges which depend on the choice of the integrable perturbation). In the case of the Lee-Yang Model that we have analysed in this paper, the KdV hierarchy and the Bullough-Dodd hierarchies exactly coincide. It is then natural to ask about GGEs with Bullough-Dodd charges inserted in them in a more general setting.
There is also the BO2 hierarchy that exists for CFTs that have a $U(1)$ current. GGEs with these charges inserted have been studied in [34, 35]. Studying their modular properties is an open question that would be interesting to explore.
@@
The modular properties of this GGE is still an open question. The first few terms in the asymptotic expansion and their modular transforms were calculated in [36, 25]. The additional charges and their thermal correlators have recently been calculated in [37, 38]. Putting these results together could allow us to find an asymptotic expression for the modular transform of (187) similar to our results in section 3. If TBA equations for the additional charges are known then we may hope to repeat the arguments of section 5 to find the full modular transform of (187).
@@ Acknowledgements
We would like to thank O. Castro-Alvaredo and A. Konechny for useful discussions during the completion of this work. We would especially like to thank G. M. T. Watts for many extensive and valuable discussions that aided in the progression of this work, reading early drafts of the paper and suggesting useful edits.
@@ Appendix A Modular forms
In this appendix we will list the relevant facts about modular forms that appear in this paper. Proofs of the following statements can be found in [39] and most of the notation will be the same.
@@ Appendix B Construction of Charges
In this appendix we will explain how to construct the charges used throughout this paper. These charges are the zero modes of quasi-primary fields on the cylinder so we will begin by explaining how we use the algorithm of Gaberdiel in [40] to map fields from the cylinder to the plane. We will then apply this map to the case of the quasi-primary field at level 6 that is linearly independent from the quasi-primary field that gives the KdV charge $I_{5}$. Finally we will discuss the charges in the Lee-Yang minimal model and give explicit expressions for all the charges used in this work.
@@ B.1 Mapping between the cylinder and plane
The algorithm for computing the $R_{n}$ is given in [40], we have listed the relevant ones for our calculations in table 5. We note that for $n$ odd and greater than 1 $R_{n}=0$ so we have not listed them in table 5. (For a general conformal map $z=f(w)$ the $R_{n}$ will be functions of $w$ and $z^{L_{0}}$ becomes $f^{\prime}(z)$.) Although we have an infinite product and the exponentials also contain infinite products these expressions can truncate to a finite one since any operator of the form $L_{n_{1}}\dots L_{n_{i}}$ with $n_{1}+\dots+n_{i}>N$ will annihilate $|\phi\rangle$.
@@ B.2 Example of a Weight 6 Field
The brackets denote normal ordering as defined in [41]. This then gives use the field $\phi^{\text{cyl}}(w)$ on the cylinder
@@
Zero modes of local operators in 2d CFTs have been discussed before in the literature. See for example [42] for a similar discussion on zero-modes of quasi-primary fields on the cylinder.
@@ B.3 Charges in the Lee-Yang Model
In this section we will present the charges relevant to the Lee-Yang model. These will include the KdV charges which have been calculated previously (see for example [3]). However we will find new simpler expressions for them by using a bases of states in the Lee-Yang theory which already has null states removed. We will also calculate the charges associated with the other quasi-primary fields in the theory. While these don’t commute with the KdV charges they still appear when we take the modular transform of a GGE as see in section 4.
There is natural basis of states in the vacuum module which avoids the null vectors in the Lee-Yang model [43], that is the vacuum module is given by
@@
Hence we have no quasi-primary states at level 4. This means we have no $I_{3}$ KdV charge as has previously been pointed out in [14].
@@
This is proportional to the state which gives the KdV charge $I_{5}$. However if we map this state to the cylinder then we get an expression for the zero mode which only contains quadratic and linear terms in the Virasoro modes rather than the usual expression for $I_{5}$ which contains terms that are cubic (see for example the expression in [3]). Hence using the representation (212) for the vacuum module leads to simpler expressions for the charges.
@@
The $I_{2n-1}$ are the KdV charges coming from a weight $2n$ quasi-primary field. They can be uniquely fixed (up to a factor) by imposing that they all commute [3]. $J_{9}$ is the zero mode of the unique quasi-primary field at level 10, it does not commute with the KdV charges. $J_{13}$ is the zero mode of the quasi-primary field at level 14 that is linearly independent to the field that gives the KdV charge $I_{13}$, it also doesn’t commute with the KdV charges.
@@ Appendix C Eigenvalues and Thermal Correlation Functions of Charges
In this appendix we explain how to compute the thermal correlation functions of the charges using the techniques of [14]. These thermal correlation functions are given by modular differential operators acting on the characters of the theory.
@@ C.2 Thermal Correlation Functions
Using the method presented in the previous section to calculate sums of eigenvalues of operators, we can find the expression of thermal correlation functions in terms of modular differential operators acting on the characters of the CFT as was done in [14].
@@
From [14] we know that the thermal correlator must be a weight 6 modular differential operator acting on the characters. We only have a linear $L_{0}$ term in (211), hence we must have a first order differential operator. Using the definition of the Serre derivative $D$ and Eisenstein series $E_{2k}$ in appendix A the only first order weight 6 differential operator we can construct is
@@
If we restrict to the Lee-Yang model we can find simpler expressions for the correlation functions compared to those given in [14] for generic CFTs. This is because the characters satisfy a second order differential equation which can be used to reduce the order of differential operator acting on the characters.
We will demonstrate this for the case of the correlator $\langle{I_{5}}\rangle$ in the Lee-Yang model. (For the rest of this section we will drop the $i$ subscript since the specific representation won’t be important.) From [14], $\langle{I_{5}}\rangle$ in the Lee-Yang model is
@@
However we also have the modular differential equation satisfied by the two characters of Lee-Yang [23]
@@ Appendix D Numerical algorithm for TBA
In this appendix we will briefly outline our approach to solving the TBA equations numerically. We do this by discretising the integrals into finite sums and then setting up iteration schemes. Our iteration scheme for finding the pseudo energy $\epsilon(\theta)$ is the same as the one used in equations (2.2) of [27] with $a=1$. We will also explicitly give the iteration scheme we used to solve for the poles $\eta$ in the excited states.